# Covering Multiple Objectives with a Small Set of Solutions Using Bayesian Optimization

### v1

@@ Abstract
In multi-objective black-box optimization, the goal is typically to find solutions that optimize a set of $T$ black-box objective functions, $f_{1},\ldots f_{T}$, simultaneously. Traditional approaches often seek a single Pareto-optimal set that balances trade-offs among all objectives. In this work, we introduce a novel problem setting that departs from this paradigm: finding a smaller set of $K$ solutions, where $K<T$, that collectively “covers” the $T$ objectives. A set of solutions is defined as “covering” if, for each objective $f_{1},\ldots f_{T}$, there is at least one good solution. A motivating example for this problem setting occurs in drug design. For example, we may have $T$ pathogens and aim to identify a set of $K<T$ antibiotics such that at least one antibiotic can be used to treat each pathogen. To address this problem, we propose Multi-Objective Coverage Bayesian Optimization (MOCOBO), a principled algorithm designed to efficiently find a covering set. We validate our approach through extensive experiments on challenging high-dimensional tasks, including applications in peptide and molecular design. Experiments demonstrate MOCOBO’s ability to find high-performing covering sets of solutions. Additionally, we show that the small sets of $K<T$ solutions found by MOCOBO can match or nearly match the performance of $T$ individually optimized solutions for the same objectives. Our results highlight MOCOBO’s potential to tackle complex multi-objective problems in domains where finding at least one high-performing solution for each objective is critical.
Bayesian optimization, coverage optimization, high-dimensional black-box optimization, multi-objective Bayesian optimization, drug design
@@ 1 Introduction
Bayesian optimization (BO) (Jones et al., 1998; Shahriari et al., 2015; Garnett, 2023) is a general framework for sample-efficient optimiziation of black-box functions. By using a probabilistic surrogate model, such as a Gaussian process, Bayesian optimization balances exploration and exploitation to identify high-performing solutions with a limited number of function evaluations. BO has been successfully applied in a wide range of domains, including hyperparameter tuning (Snoek et al., 2012b; Turner et al., 2021), A/B testing (Letham et al., 2019), chemical engineering (Hernández-Lobato et al., 2017), drug discovery (Negoescu et al., 2011), and more.
In multi-objective optimization the typical assumption is that objectives inherently trade off with one another, motivating the search for Pareto-optimal solutions that balance these trade-offs. Multi-objective optimization methods focus on identifying diverse (in objective value) sets of solutions along the Pareto front to capture these trade-offs (Hernández-Lobato et al., 2015; Belakaria et al., 2019; Turchetta et al., 2019; Konakovic Lukovic et al., 2020; Daulton et al., 2021; Stanton et al., 2022; Belakaria et al., 2022). Indeed, without such trade-offs, optimization would trivialize into simply optimizing all of the objectives.
@@
Figure 2: Diagram of example “coverage optimization” problem. MOCOBO seeks a set of $K=4$ peptides to cover $T=11$ target pathogenic bacteria. Arrows from peptides to bacteria indicate which peptides successfully inhibit (or “cover”) each bacteria in the target set. We provide results for this problem setting in Section 4.
To formalize this, we propose the “coverage optimization” problem: the task of covering a set of $T$ objectives with $K$ solutions, where coverage means that for each objective $f_{1},\dots,f_{T}$, there exists at least one high-performing solution among the $K$. In the context of drug design, this translates to identifying $K$ drugs such that each of the $T$ pathogens is effectively addressed by at least one drug in the set. This problem setting diverges from traditional multi-objective optimization by prioritizing collective coverage over balancing trade-offs within individual solutions. In Figure 2, we provide a diagram of one such “coverage optimization” problem considered in this paper.
To address this novel problem setting, we introduce and experimentally validate Multi-Objective Coverage Bayesian Optimization (MOCOBO), a new algorithm designed to solve the “coverage optimization” problem within the Bayesian optimization framework. MOCOBO efficiently identifies a small set of $K$ solutions that collectively cover $T$ objectives. Unlike traditional multi-objective methods that seek to balance trade-offs across all objectives for every solution, MOCOBO focuses on identifying specialized sets of solutions to meet practical coverage requirements.
@@
    We introduce the problem setting of finding a “covering set” of $K<T$ high-performing solutions such that each relevant objective $f_{t}\in\left\{f_{1},\ldots f_{T}\right\}$ is optimized by at least one solution. This is the first work we are aware of to consider this problem setting.
@@
    We propose a local Bayesian optimization algorithm, MOCOBO, for this problem setting, that extends to high dimensional objectives, structured input domains, and large numbers of function evaluations.
@@
    We validate MOCOBO through experiments on challenging, high-dimensional optimization tasks including structured drug discovery over molecules and peptides, demonstrating its ability to consistently outperform state-of-the-art Bayesian optimization methods in identifying high-performing covering sets of solutions.
@@
    We provide empirical results demonstrating that the small sets $K<T$ solutions found by MOCOBO consistently nearly match the performance of a larger set of $T$ solutions individually optimized for each task.
@@
    We demonstrate the potential of MOCOBO for drug discovery using an in vitro experiment showing it produces potent antimicrobial peptides that cover 9/11 drug resistant or otherwise challenging to kill pathogens, with moderate activity on 1/11.
@@ 2 Background and Related Work
BO (Močkus, 1975; Snoek et al., 2012a) is a methodology for sample-efficient black-box optimization. BO operates iteratively, where each cycle involves training a probabilistic surrogate model, typically a Gaussian process (GP) (Rasmussen, 2003), on data collected from the black-box objective function. BO then employs the predictive posterior $p(y\mid\mathbf{x},D)$ of the surrogate model to compute an acquisition function that gives a policy for determining the next candidate(s) to evaluate, balancing exploration and exploitation.
@@
Recent advances in MOBO have introduced numerous methods aimed at identifying diverse solutions along the Pareto front to balance trade-offs among multiple objectives (Hernández-Lobato et al., 2015; Belakaria et al., 2019; Turchetta et al., 2019; Konakovic Lukovic et al., 2020; Daulton et al., 2021; Stanton et al., 2022; Belakaria et al., 2022). These approaches, however, are not well-suited to “coverage optimization”, where the goal is to find a small set of $K<T$ solutions that collectively address all $T$ objectives. Despite this, we compare our proposed method to MORBO, developed by Daulton et al. (2021), which is regarded as a state-of-the-art method for high-dimensional MOBO tasks.
@@
BO has recently been used for optimizing structured search spaces like molecules or amino acid sequences by employing latent space Bayesian optimization, which utilizes a variational autoencoder (VAE) to convert structured inputs into a continuous latent space for BO (Gómez-Bombarelli et al., 2018; Eissman et al., 2018; Tripp et al., 2020; Grosnit et al., 2021; Siivola et al., 2021; Jin et al., 2018; Stanton et al., 2022). The VAE encoder, $\Phi(\mathbf{x})$, transforms structured inputs $\mathbf{x}$ into continuous latent vectors $\mathbf{z}$, enabling BO to operate in the latent space. Candidate latent vectors are decoded using the VAE decoder, $\Gamma(\mathbf{z})$, to generate structured outputs. Maus et al. (2022) extended TuRBO (Eriksson et al., 2019) for latent space optimization by jointly training the surrogate model and VAE using variational inference.
@@
Local Bayesian optimization (Eriksson et al., 2019), including TuRBO-$M$, is one recent successful approach to high-dimensional BO. TuRBO-$M$ uses $M$ parallel local optimization runs, each maintaining its own dataset $\mathcal{D}_{i}$ and surrogate model. Each local optimizer proposes candidates only within a hyper-rectangular trust region $\mathcal{T}_{i}$. The trust region $\mathcal{T}_{i}$ is a rectangular area of the input space $\mathcal{X}$, centered on the incumbent $\mathbf{x}^{+}_{i}$. Its side length $\ell_{i}$ is bound by $[\ell_{min},\ell_{max}]$. If an optimizer improves the incumbent for $\rho_{succ}$ consecutive iterations, $\ell_{i}$ expands to $\min(2\ell_{i},\ell_{max})$. If not improved in $\rho_{fail}$ iterations, $\ell_{i}$ is halved. Optimizers are restarted if $\ell_{i}$ drops below $\ell_{min}$. While TuRBO-$M$ is not directly applicable to our problem, we follow methods like MORBO (Daulton et al., 2021) and ROBOT (Maus et al., 2023) in adapting the use of $M$ coordinated trust regions ($K$ in our case) to a different problem setting.
@@ 3 Methods
We consider the task of finding a set of $K<T$ solutions $S^{*}=\left\{\mathbf{x}^{*}_{1},\ldots,\mathbf{x}^{*}_{K}\right\}$ that “covers” a set of $T$ objectives $f_{1},\ldots,f_{T}$. Here, the set $S^{*}$ “covers” the $T$ objectives if, for each objective $f_{t}=f_{1},\ldots,f_{T}$, there is at least one solution $\mathbf{x}^{*}_{k}\in S^{*}$ for which $f_{t}$ is well optimized by $\mathbf{x}^{*}_{k}$. We evaluate how well a set of $K$ points $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}$ “covers” the $T$ objectives using following “coverage score” $c$:
@@ 3.1 Multi-Objective Coverage Bayesian Optimization (MOCOBO)
In this section, we propose MOCOBO - an algorithm which extends Bayesian optimization to the problem setting above.
In order to find a set of $K$ solutions, MOCOBO maintains $K$ simultaneous local optimization runs using $K$ individual trust regions. Each local run $k\in\left\{1,\ldots,K\right\}$ aims to find a single solution $\mathbf{x}^{*}_{k}$, which together form the desired set $S^{*}$. As in the original TuRBO paper, trust regions are rectangular regions of the search space $\mathcal{T}_{k}\subseteq\mathcal{X}$ defined solely by their size and center point.
On each step of optimization $s$, we use our current set of all data evaluated so far $D_{s}=\left\{(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{n},\mathbf{y}_{n})\right\}$ to define the best covering set $S^{*}_{D_{s}}=\left\{\mathbf{x}^{*(s)}_{1},\ldots,\mathbf{x}^{*(s)}_{K}\in D_{s}\right\}$ found so far. Here $\mathbf{y}_{i}=(f_{1}(\mathbf{x}_{i}),\ldots,f_{T}(\mathbf{x}_{i}))$ and $n$ is the number of data points evaluated so far at step $s$. Following Equation 2, we define $S^{*}_{D_{s}}$ as follows:
$S^{*}_{D_{s}}$ $=\operatorname*{arg\,max}_{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}\subseteq D_{s}}c(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\})$ (3)
$=\operatorname*{arg\,max}_{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}\subseteq D_{s}}\sum_{t=1}^{T}\max_{k=1}^{K}f_{t}(\mathbf{x}_{k})$ (4)
@@
Since we select and evaluate $q$ candidates from each of the $K$ local optimization runs on each step of optimization, $n=s\times K\times q$.
@@ 3.1.1 Candidate Selection with Expected Coverage Improvement (ECI)
$ECI(\mathbf{x})$ $=\mathbb{E}_{p(y|\mathbf{x},D)}[\max(0,c(S^{*}_{D_{s}\cup\left\{(\mathbf{x},\mathbf{y})\right\}})-c(S^{*}_{D_{s}}))]$ (5)
@@
We estimate ECI using a Monte Carlo (MC) approximation. To select a single candidate $\hat{\mathbf{x}}_{k}$ from $\mathcal{T}_{k}$, we sample $m$ points $P_{k}=\left\{\mathbf{p}_{k1},\mathbf{p}_{k2},...,\mathbf{p}_{km}\right\}$ from $\mathcal{T}_{k}$. For each sampled point $\mathbf{p}_{kj}$, we sample a realization $\hat{\mathbf{y}}_{kj}=(\hat{f}_{1}(\mathbf{p}_{kj}),\ldots,\hat{f}_{T}(\mathbf{p}_{kj}))$ from the GP surrogate model posterior. We leverage these samples to compute an MC approximation to the ECI of each $\mathbf{p}_{kj}$:
$CI(\mathbf{p}_{kj})$ $=\max(0,c(S^{*}_{D_{s}\cup\left\{(\mathbf{p}_{kj},\hat{\mathbf{y}}_{kj})\right\}}-c(S^{*}_{D_{s}})))$ (6)
Here, $c(S^{*}_{D_{s}\cup\left\{(\mathbf{p}_{kj},\hat{\mathbf{y}}_{kj})\right\}})$ is the approximation of the coverage score of the new best covering set if we choose to evaluate candidate $\mathbf{p}_{kj}$, assuming the candidate point will have the sampled objective values $\hat{\mathbf{y}}_{kj}$. We select and evaluate the candidate $\mathbf{p}_{kj}$ with largest expected coverage improvement.

### v2

# Multi-Objective Coverage Bayesian Optimization (MOCOBO)

@@ Abstract
In multi-objective black-box optimization, the goal is typically to find solutions that optimize a set of $T$ black-box objective functions, $f_{1},\ldots f_{T}$, simultaneously. Traditional approaches often seek a single Pareto-optimal set that balances trade-offs among all objectives. In this work, we consider a problem setting that departs from this paradigm: finding a small set of $K<T$ solutions, that collectively “covers" the $T$ objectives. A set of solutions is defined as “covering" if, for each objective $f_{1},\ldots f_{T}$, there is at least one good solution. A motivating example for this problem setting occurs in drug design. For example, we may have $T$ pathogens and aim to identify a set of $K<T$ antibiotics such that at least one antibiotic can be used to treat each pathogen. To address this problem, we propose Multi-Objective Coverage Bayesian Optimization (MOCOBO), a principled algorithm designed to efficiently find a covering set. We validate our approach through experiments on challenging high-dimensional tasks, including applications in peptide and molecular design, where MOCOBO is shown to find high-performing covering sets of solutions. The results show that the coverage of the $K<T$ solutions found by MOCOBO matches or nearly matches the coverage of $T$ solutions obtained by optimizing each objective individually. Furthermore, in in vitro experiments, the peptides found by MOCOBO exhibited high potency against drug-resistant pathogens, further demonstrating the potential of MOCOBO for drug discovery.
@@ 1 Introduction
Bayesian optimization (BO; 31, 54, 21) is a general framework for sample-efficient optimiziation of black-box functions. By using a probabilistic surrogate model, such as a Gaussian process, Bayesian optimization balances exploration and exploitation to identify high-performing solutions with a limited number of function evaluations. BO has been successfully applied in a wide range of domains, including hyperparameter tuning [56, 64], A/B testing [37], chemical engineering [26], drug discovery [47], and more.
In multi-objective optimization the typical assumption is that objectives inherently trade off with one another, motivating the search for Pareto-optimal solutions that balance these trade-offs. Multi-objective Bayesian optimization methods focus on identifying diverse (in objective value) sets of solutions along the Pareto front to capture these trade-offs [27, 3, 63, 33, 11, 57, 4]. Indeed, without such trade-offs, optimization would trivialize into simply optimizing all of the objectives.
@@
To formalize this, we consider the “coverage optimization" problem: the task of covering a set of $T$ objectives with $K$ solutions, where coverage means that for each objective $f_{1},\dots,f_{T}$, there exists at least one high-performing solution among the $K$. In the context of drug design, this translates to identifying $K$ drugs such that each of the $T$ pathogens is effectively addressed by at least one drug in the set. This problem setting diverges from traditional multi-objective optimization by prioritizing collective coverage over balancing trade-offs within individual solutions.
To address the coverage problem setting, we introduce and experimentally validate Multi-Objective Coverage Bayesian Optimization (MOCOBO), a new algorithm designed to solve the “coverage optimization" problem within the Bayesian optimization framework. MOCOBO efficiently identifies a small set of $K$ solutions that collectively cover $T$ objectives. Unlike traditional multi-objective Bayesian optimization methods that seek to balance trade-offs across all objectives for every solution, MOCOBO focuses on identifying specialized sets of solutions to meet practical coverage requirements.
@@
    We consider the problem setting of finding a “covering set" of $K<T$ high-performing solutions such that each relevant objective $f_{t}\in\left\{f_{1},\ldots f_{T}\right\}$ is optimized by at least one solution. As far as we are aware, this is the first work in the Bayesian optimization (BO) literature to consider the coverage optimization problem setting.
@@
    We propose a local BO algorithm, MOCOBO, for this problem setting, that extends to high dimensional objectives, structured input domains, and large numbers of function evaluations.
@@
    We experimentally validate MOCOBO on challenging, high-dimensional optimization tasks including structured drug discovery over molecules and peptides, demonstrating its ability to consistently outperform state-of-the-art BO methods in identifying high-performing covering sets of solutions.
@@
    We demonstrate empirically that the small sets of $K<T$ solutions found by MOCOBO consistently nearly match the performance of a larger set of $T$ solutions individually optimized for each task.
@@
    We demonstrate the potential of MOCOBO for drug discovery using an in vitro experiment showing it produces potent antimicrobial peptides that cover 9 out of 11 drug resistant or otherwise challenging to kill pathogens, with moderate activity on 1 out of 11.
@@ 2 Background and related work
BO is a method for sample-efficient black-box optimization. It operates iteratively, where each cycle involves training a probabilistic surrogate model, typically a Gaussian process (GP; 50), on data collected from the black-box objective function. BO then employs the predictive posterior $p(y\mid\mathbf{x},D)$ of the surrogate model to compute an acquisition function that gives a policy for choosing the next candidate(s) to evaluate, balancing exploration and exploitation.
@@
Recent advances in MOBO have introduced numerous methods aimed at identifying diverse solutions along the Pareto front to balance trade-offs among multiple objectives [27, 3, 63, 33, 11, 57, 4]. These approaches, however, are not well-suited to “coverage optimization”, where the goal is to find a small set of $K<T$ solutions that collectively address all $T$ objectives. Despite this, we compare our proposed method to MORBO, developed by Daulton et al. [11], which is regarded as a state-of-the-art method for high-dimensional MOBO tasks.
@@
BO has recently been used for optimizing structured search spaces like molecules or amino acid sequences by employing latent space Bayesian optimization, which utilizes a variational autoencoder (VAE; 32, 52) to convert structured inputs into a continuous latent space for BO [23, 15, 61, 24, 55, 30, 57]. The VAE encoder, $\Phi(\mathbf{x})$, transforms structured inputs $\mathbf{x}$ into continuous latent vectors $\mathbf{z}$, enabling BO to operate in the latent space. Candidate latent vectors are decoded using the VAE decoder, $\Gamma(\mathbf{z})$, to generate structured outputs. Maus et al. [44] extended TuRBO [17] for latent space optimization by jointly training the surrogate model and VAE using variational inference.
@@
Local Bayesian optimization [17], including TuRBO-$M$, is one recent successful approach to high-dimensional BO. TuRBO-$M$ uses $M$ parallel local optimization runs, each maintaining its own dataset $\mathcal{D}_{i}$ and surrogate model. Each local optimizer proposes candidates only within a hyper-rectangular trust region $\mathcal{T}_{i}$. The trust region $\mathcal{T}_{i}$ is a rectangular area of the input space $\mathcal{X}$, centered on the incumbent $\mathbf{x}^{+}_{i}$. Its side length $\ell_{i}$ is bound by $[\ell_{min},\ell_{max}]$. If an optimizer improves the incumbent for $\rho_{succ}$ consecutive iterations, $\ell_{i}$ expands to $\min(2\ell_{i},\ell_{max})$. If not improved in $\rho_{fail}$ iterations, $\ell_{i}$ is halved. Optimizers are restarted if $\ell_{i}$ drops below $\ell_{min}$. While TuRBO-$M$ is not directly applicable to our problem, we follow methods like MORBO [11] and ROBOT [45] in adapting the use of $M$ coordinated trust regions ($K$ in our case) to a different problem setting.
Prior work on coverage optimization and related problem settings.
To our knowledge, the coverage optimization problem setting considered in this paper has not been previously explored in the Bayesian optimization literature. However, similar problem settings have been studied in the context of gradient-based multi-objective optimization. In particular, Ding et al. [13], Li et al. [39], and Lin et al. [40] have investigated related problems in this domain. Ding et al. [13] demonstrate that the coverage optimization problem generalizes various clustering methods, including k-means clustering, and propose an algorithm that extends k-means++ [1] and Lloyd’s algorithm [43] to efficiently solve the coverage problem using gradient descent. Li et al. [39] introduce Many-objective Multi-solution Transport (MosT), a novel framework that scales multi-objective gradient-based optimization to a large number of objectives. MosT uses an optimal transport to establish a balanced assignment between objectives and solutions, allowing each solution to focus on a specific subset of objectives and ensuring that the collective set of solutions covers all objectives. Lin et al. [40] introduce a novel Tchebycheff set (TCH-Set) scalarization approach for covering multiple conflicting objectives using a small set of collaborative solutions. They further propose a smooth Tchebycheff set (STCH-Set) scalarization method to handle non-smoothness in TCH-Set scalarization. While these three papers address related coverage optimization problems, their proposed methods rely on differentiable objective functions, making them unsuitable for the black-box optimization problems considered in this paper. Liu et al. [42] were the first to consider the coverage problem in the setting of black-box optimization. They propose a clustering-based swarm optimization algorithm (CluSO) for the coverage problem in the setting of multi-objective black-box optimization. In Section 4, we provide empirical results demonstrating that our Bayesian optimization approach (MOCOBO) consistently outperforms CluSO by a large margin across all tasks considered.
@@ 3 Methods
We consider the task of finding a set of $K<T$ solutions $S^{*}=\left\{\mathbf{x}^{*}_{1},\ldots,\mathbf{x}^{*}_{K}\right\}$ that “covers" $T$ objectives $f_{1},\ldots,f_{T}$. Here, the set $S^{*}$ “covers" the $T$ objectives if, for each objective $f_{t}=f_{1},\ldots,f_{T}$, there is at least one solution $\mathbf{x}^{*}_{k}\in S^{*}$ for which $f_{t}$ is well optimized by $\mathbf{x}^{*}_{k}$. We evaluate how well a set of $K$ points $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}$ “covers" the $T$ objectives using the “coverage score" $c$:
@@ 3.1 Multi-Objective Coverage Bayesian Optimization (MOCOBO)
In this section, we propose MOCOBO - an algorithm which extends Bayesian optimization to the problem setting above. On each step of optimization $s$, we use our current set of all data evaluated so far $D_{s}=\left\{(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{n},\mathbf{y}_{n})\right\}$ to define the best covering set $S^{*}_{D_{s}}=\left\{\mathbf{x}^{*(s)}_{1},\ldots,\mathbf{x}^{*(s)}_{K}\in D_{s}\right\}$ found so far. Here $\mathbf{y}_{i}=(f_{1}(\mathbf{x}_{i}),\ldots,f_{T}(\mathbf{x}_{i}))$ and $n$ is the number of data points evaluated so far at step $s$. Following Equation 2, we define $S^{*}_{D_{s}}$ as follows:
$S^{*}_{D_{s}}$ $=\operatorname*{arg\,max}_{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}\subseteq D_{s}}c(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\})$ $=\operatorname*{arg\,max}_{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}\subseteq D_{s}}\sum_{t=1}^{T}\max_{k=1}^{K}f_{t}(\mathbf{x}_{k})$ (3)
@@ 3.1.1 Candidate selection with Expected Coverage Improvement (ECI)
$\mathrm{ECI}(\mathbf{x})$ $=\mathbb{E}_{p(y|\mathbf{x},D)}[\max(0,c(S^{*}_{D_{s}\cup\left\{(\mathbf{x},\mathbf{y})\right\}})-c(S^{*}_{D_{s}}))]$ (4)
@@
We estimate ECI using a Monte Carlo (MC) approximation. To select a single candidate $\hat{\mathbf{x}}$, we sample $m$ points $P=\left\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{m}\right\}$. For each sampled point $\mathbf{p}_{j}$, we sample a realization $\hat{\mathbf{y}}_{j}=(\hat{f}_{1}(\mathbf{p}_{j}),\ldots,\hat{f}_{T}(\mathbf{p}_{j}))$ from the GP surrogate model posterior. We leverage these samples to compute an MC approximation to the ECI of each $\mathbf{p}_{j}$:
$\mathrm{CI}(\mathbf{p}_{j})$ $=\max(0,c(S^{*}_{D_{s}\cup\left\{(\mathbf{p}_{j},\hat{\mathbf{y}}_{j})\right\}}-c(S^{*}_{D_{s}})))$ (5)
Here, $c(S^{*}_{D_{s}\cup\left\{(\mathbf{p}_{j},\hat{\mathbf{y}}_{j})\right\}})$ is the approximation of the coverage score of the new best covering set if we choose to evaluate candidate $\mathbf{p}_{j}$, assuming the candidate point will have the sampled objective values $\hat{\mathbf{y}}_{j}$. We select and evaluate the candidate $\mathbf{p}_{j}$ with largest expected coverage improvement.